OpenStudy (anonymous):

10*9*8*7*6*5*4*3*2*1 = 10! Can this be true?! Why or why not?

OpenStudy (debbieg):

Do you know the definition of a factorial? n!=n(n-1)(n-2)(n-3)*...*1

OpenStudy (anonymous):

The expression "x!" means x(x-1)(x-2)(x-3)...1 So 10! means 10(10-1)(10-2)(10-3)(10-4)(10-5)(10-6)(10-7)(10-8)(10-9)= 10*9*8*7*6*5*4*3*2*1=10!, which means it is correct.

OpenStudy (anonymous):

thanks for the help

OpenStudy (anonymous):

False....

OpenStudy (debbieg):

How so, @whalexnuker ? @DUDE..IM..A..DUCK , yes, that's correct.

OpenStudy (anonymous):

i know thanks again :D

OpenStudy (debbieg):

you're welcome. :)

OpenStudy (amistre64):

as long as you are able to define the operation/notation ; just about anything can be "true" when applied to a proper context.